OpenStudy (anonymous):

81.5°F=_____Celcius a.13.3 b.27.5 c.178.7 d.204.3

OpenStudy (anonymous):

b

OpenStudy (anonymous):

B 27.5*1.8 + 32 = 81.5

OpenStudy (anonymous):

Conversione da, a, Formula. celsius, fahrenheit, °F = (9/5) × °C + 32. fahrenheit, celsius, °C = (5/9) × (°F - 32). celsius, kelvin, K = °C + 273,15 you can use these formulae